OpenStudy (kweenlexis):

Which ordered pairs in the form (x, y) are solutions to the equation 7x−5y=28? Select each correct answer. (−1, −7) (−6, −14) (7, 9) (4, 10)

OpenStudy (mhchen):

Did you try to plug them in? You can convert it into 5y = 7x - 28 y = (7x/5) - (28/5) And check.

OpenStudy (dallascowboys88):

D......

OpenStudy (kweenlexis):

@mhchen ??

OpenStudy (mhchen):

So the ordered pairs are (x, y) For example, (-1, -7) --> -7 = (7(-1)/5) -(28/5) <-- Check to see if that equals. If it doens't, move onto the next pair.

OpenStudy (kweenlexis):

idk hoow to do it...

OpenStudy (mhchen):

Just plug it in. Like for the first one \[y = \frac{ 7x }{ 5 } - \frac{ 28 }{ 5 }\] For (-1, -7) \ [-7 = \frac{ 7*(-1) }{ 5 } - \frac{ 28 }{ ?5 }\] And then you see if the above equation makes sense. If it DOES, then that's your answer. If it DOESN'T, then just move onto the next one.

OpenStudy (kweenlexis):

ok

OpenStudy (kweenlexis):

I'm lowkey still lost

OpenStudy (mhchen):

Ohai Lowkey. So if you look at our equation, it has an x and a y. Right?

OpenStudy (kweenlexis):

Yeah

OpenStudy (mhchen):

and for the 4 ordered pairs you have, it's ( x-vaue , y- value)

OpenStudy (mhchen):

So you put the x-value into the x part of the equation, and the y-value into the y part of the equation.

OpenStudy (kweenlexis):

i dont understand sooo im jusst guess the answer

OpenStudy (mhchen):

And our equation is basically the line on a graph, if the points (which is the ordered pairs) is on the line, then the equation should equal.

OpenStudy (mhchen):

If you're gonna guess, guess (A)

OpenStudy (kweenlexis):

they say its 2 answers tho

OpenStudy (mhchen):

Ah crap, so there's another one.

OpenStudy (kweenlexis):

yeah

OpenStudy (mhchen):

You can just do the old fashion way of plugging x and y values into the normal equation. 7(-1) - 5(-7) = 28 7(-6) - 5(-14) = 28 7(7) - 5(9) = 28 7(4) - 5(10) = 28

OpenStudy (mhchen):

So the first equation works, -7 + 25 = 28 The 2nd equation is -42 + 70 = 28 So it's (A) and (B)

OpenStudy (kweenlexis):

wait do you doo...7*(-1 1st on A?

OpenStudy (mhchen):

Yeah, sorry i g2g rn.

OpenStudy (kweenlexis):

mhm